# Inductive Gradient Adjustment for Spectral bias in implicit neural representations

### v1

@@ Appendix C Empirical Analysis on Simple Function Approximation
Appendix E Experimental details and per-scene results of 3D shape representation
In this section, we firstly introduce the detailed training strategy of the 3D shape representation experiment. Then we provide the intersection over union (IOU) metric value of each 3D shape and give more visualization.
Following previous works (Saragadam et al., 2023; Shi et al., 2024a; Cai et al., 2024), we train all models for 200 epochs with a learning rate decay exponentially to $0.1$ of initial rates. We set the initial learning rate as $32e-$ 3 for ReLU and PE; For SIREN, we set the initial learning rate as $45e-$ 4. FR, BN and our IGA method are adopted the same learning rate with baseline models. We set $end=7$ for ReLU and PE; For SIREN, we set $end=14$. The per-scene results are reported in Table 9. It can be observed that our IGA method achieves improvements in IOU metrics across all objects, and achieves the best performance in average metrics of all objects. We visualize the Thai and Lucy objects in Fig. 11.
Table 9: Intersection over Union (IOU) of 3D shape representation by different methods. The detailed settings can be found in Sec. 6.2 and Appendix E.
@@ Appendix E Experimental details and per-scene results of 3D shape representation
Figure 11: Visualization of 3D shape representation results and IOU values on the Thai (top) and the Lucy (bottom). ReLU+IGA denotes that the MLPs with ReLU activation function are optimized by our IGA method. Details can be found in Sec. 6.2.
Appendix F Experimental details and per-scene results of learning neural radiance fields
In this section, we firstly introduce some training details. Then we provide metrics of each scenario and give more visualization. As previously discussed in Experiment 6.3, we apply our method on the original NeRF model (Mildenhall et al., 2021). The original NeRF adopts two models: “coarse” and “fine” models. We adjust the gradients of the “fine” model as it captures more high-frequency details and remain gradients of the “coarse” model. The “NeRF-pytorch” codebase (Yen-Chen, 2020) is used and we follow its default settings. Besides, we also compare to previous training dynamics methods, i.e., FR and BN. Their hyperparameters follow their publicly codes. Generally, we set the $end=25$. For more complex scenes, such as the branches and leaves of the ficus, we set the $end=30$ to achieve better learning of high frequencies. We report per-scene results in Table 10 and visualize several scenes in Fig. 12.
Table 10: Per-scene results of learning 5D neural radiance fields by different methods. Details can be found in Sec. 6.3 and Appendix F.
@@ Appendix F Experimental details and per-scene results of learning neural radiance fields
Figure 12: Visual examples of novel view synthesis results of NeRF and NeRF+IGA. Details can be found in Sec. 6.3 and Appendix F.

### v2

@@ Appendix I Experimental details and per-scene results of Learning Neural Radiance Fields
Table 10: Multi-Scale Structural Similarity Index Measure (MS-SSIM $\uparrow$) of 2D color image approximation results by different methods. The detailed settings can be found in Sec. 5.1.
ReLU 0.5270 0.7839 0.8414 0.7741 0.4697 0.6534 0.6646 0.5031 0.6521
ReLU+FR 0.5454 0.7973 0.8621 0.7964 0.5164 0.6679 0.7024 0.5523 0.6800
ReLU+BN 0.5837 0.8075 0.8581 0.8047 0.5943 0.6929 0.7331 0.5976 0.7090
ReLU+IGA 0.6205 0.8364 0.8947 0.8249 0.6208 0.7290 0.7591 0.6211 0.7383
PE 0.9356 0.9518 0.9668 0.9489 0.9382 0.9331 0.9744 0.9242 0.9466
PE+FR 0.9493 0.9494 0.9713 0.9559 0.9566 0.9470 0.9755 0.9464 0.9564
PE+BN 0.9491 0.9507 0.9615 0.9560 0.9561 0.9481 0.9726 0.9491 0.9554
PE+IGA 0.9705 0.9720 0.9880 0.9788 0.9713 0.9692 0.9895 0.9622 0.9752
SIREN 0.9808 0.9817 0.9851 0.9829 0.9845 0.9714 0.9918 0.9761 0.9818
SIREN+FR 0.9797 0.9788 0.9877 0.9819 0.9810 0.9781 0.9926 0.9766 0.9820
SIREN+IGA 0.9815 0.9851 0.9895 0.9873 0.9860 0.9803 0.9921 0.9759 0.9847
@@
Table 12: Intersection over Union (IOU) of 3D shape representation by different methods. The detailed settings can be found in Sec. 5.2 and Appendix H.
@@
Table 13: Chamfer Distance of 3D shape representation by different methods. The detailed settings can be found in Sec. 5.2 and Appendix H.
Method Thai Armadillo Dragon Bun Lucy Average
ReLU 5.6726e-06 5.5805e-06 6.5588e-06 8.0611e-06 3.8048e-06 5.9356e-06
ReLU+FR 5.2954e-06 5.6664e-06 6.5062e-06 8.2686e-06 3.5963e-06 5.8666e-06
ReLU+BN 5.3489e-06 5.4240e-06 5.8546e-06 7.8738e-06 3.5496e-06 5.6102e-06
ReLU+IGA 4.7463e-06 5.4097e-06 5.7295e-06 8.2330e-06 3.3144e-06 5.4866e-06
PE 4.2759e-06 5.2162e-06 5.3788e-06 7.7409e-06 3.0036e-06 5.1231e-06
PE+FR 4.2272e-06 5.2246e-06 5.3514e-06 7.7974e-06 2.9542e-06 5.1109e-06
PE+BN 4.2463e-06 5.2030e-06 5.4122e-06 7.7698e-06 2.9493e-06 5.1161e-06
PE+IGA 4.2551e-06 5.1524e-06 5.4319e-06 7.7695e-06 2.9288e-06 5.1076e-06
SIREN 4.2920e-06 7.8573e-06 5.3923e-06 7.9121e-06 2.9870e-06 5.6881e-06
SIREN+FR 4.3588e-06 5.1830e-06 5.4719e-06 7.7805e-06 3.0217e-06 5.1632e-06
SIREN+IGA 4.3520e-06 5.2203e-06 5.3757e-06 7.8457e-06 2.9934e-06 5.1574e-06
Table 14: Per-scene results of learning 5D neural radiance fields by different methods. Details can be found in Sec. 5.3 and Appendix I.
@@
Figure 13: Visual examples of novel view synthesis results of NeRF and NeRF+IGA. Details can be found in Sec. 5.3 and Appendix I.